OpenStudy (anonymous):

If x^2 - 6x + 8 = 0, then x - 4 = 0 or x - 2 = 0. true or false?

OpenStudy (raden):

actually, it is the correct statement. because x^2-6x+8=0 or (x-2)(x-4)=0, for zeroes, it will be x - 4 = 0 or x - 2 = 0

OpenStudy (raden):

so, true :)